A pharmaceutical company receives large shipments of ibuprofen tablets and uses this acceptance sampling plan: randomly select and test 29 tablets, then accept the whole batch if there is at most one that doesn't meet the required specifications. If a particular shipment of thousands of ibuprofen tablets actually has a 12% rate of defects, what is the probability that this whole shipment will be accepted?
Answer #1

Solution:

No. Of samples =29

P(X<=1)=?

P(success) =0.12

P(fail) =0.88

P(x<=1) = p(x=1) +p(x=0)

=(29C1 * 0.12^1 * 0.88^28)+ ( 29C0 * 0.12^0 * 0.88^29)

= 29*0.12* 0.0278 + 1 * 1 *0.0245

= 0.0967 + 0.0245 = 0.1212 = 12.12%